# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v1

@@ B.5 Proof of Corollary
These are common assumptions (see e.g., Buhlmann and Yu, 2002). The first three conditions are straightforward. The tail condition (i.e., Assumption (B.3)-(d)) rules out settings in which the density of $\tau_{i}$ concentrates at an infinitely large value. Informally, this implies that the treatment effect cannot explode to be infinitely large, which is reasonable in many substantive contexts of interest. The final condition (i.e., Assumption (B.3)-(e)) which we refer to as a signal condition, rules out settings in which there is no amount of variation in $\tau_{i}$ that can be explained by the covariates $X_{i}$ (i.e., the underlying data generating process is pure noise). In other words, the covariates we have collected must have some degree of explanatory power. This rules out the setting considered in Cattaneo et al. (2022).
@@
Under Assumption (B.3) and squared loss, we apply Buhlmann and Yu (2002), Theorem 3.1 to show that as $n\to\infty$,
@@
where $B(t)$ is a two-sided Brownian motion, originating from zero, and $V=-dF_{X}(s)f^{\prime}(s)\neq 0$. To derive the variance of $W_{\sigma^{2},s}$, we can exploit the fact that $W_{\sigma^{2},s}$ follows Chernoff’s distribution (Groeneboom, 1989).
@@
where $B(t)$ is a standard Brownian motion, and $c>0$. Then, from Groeneboom (1989) (Corollary 3.3), the density of $Z$ is given as $dF_{Z}(t)=g_{\gamma}(t)g_{\gamma}(-t),$ and $g_{\gamma}(t)$ is defined as:
@@
where $c_{1}>a_{1}$, and $a_{1}$ is the largest zero of the Airy function Ai. We can apply Janson (2013) (Theorem 1.1) to obtain a closed form solution for the variance of $Z_{\gamma}$:
@@ B.7 Proof of Lemmas
$~{}~{}~{}~{}~{}~{}~{}~{}~{}\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}-\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},s^{(a)})\right\}+\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},\hat{s}^{(a)}_{n})\right\}<0\Big{]}$
@@
$=\Pr\left(\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}-\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]+\delta_{ba}<0\right)$
$=\Pr\left(\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]-\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right]>\delta_{ba}\right)$
$\leq\Pr\left(\left|\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]-\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}_{i};X^{(b)},s^{(b)})\right\}\right]\right|>\delta_{ba}\right)$
@@
$\leq\frac{1}{\delta_{ba}}\mathbb{E}\left(\left|\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]-\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},s^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right]\right|\right)$
$\leq\frac{1}{\delta_{ba}}\left\{\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right|\right]+\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right|\right]\right\}$
@@ Appendix C Method Implementation Details
In this section, we describe the hyperparameters and implementation details for all methods used in the simulations and case study. All methods and code were implemented using R (R Core Team, 2024). For ease of reproducibility and communication of the simulation results, the simulation study was conducted using the simChef R package (Duncan et al., 2024).
@@
We considered causal distillation trees with two different teacher models: causal forest (Wager and Athey, 2018) and R-learner with boosting (rboost) (Nie and Wager, 2021). These teacher models were implemented using grf::causal_forest() (Tibshirani et al., 2024) and rlearner::rboost() (Nie et al., 2021), respectively, using their default hyperparameter settings.
For the student model, we fit a CART decision tree (Breiman et al., 1984) using rpart::rpart() (Therneau and Atkinson, 2023), with the default settings. The CART was then pruned using the standard post-pruning procedure (i.e., choosing the complexity parameter $\alpha$ which minimizes the cross-validation error), as described in Section 3.1. Note that this post-pruning step does not drastically affect the subgroup estimation results, as shown in Figure A4.
To perform honest estimation of the subgroup ATEs, we set $\pi_{train}=0.70$ for splitting the data into a training and hold-out estimation set. To estimate the individual-level treatment effects $\hat{\tau}(X_{i})$, we used the out-of-bag sample estimates (without repeated cross-fitting) in the Distilled Causal Forest and the repeated cross-fitting procedure with $R=50$ repeats in the Distilled Rboost. We also examined other choices of $R$ in Figure A7. From these simulations, we found that $R=50$ provided a good balance between the subgroup estimation accuracy and computational burden. However, more generally, the subgroup estimation results from Distilled Rboost were robust as long as $R$ was sufficiently large enough.
@@
We implemented the virtual twins algorithm (Foster et al., 2011) using the ranger::ranger() (Wright and Ziegler, 2017) implementation of random forests. Like in CDT, we set aside a hold-out estimation set to perform honest estimation of the subgroup average treatment effects and used $\pi_{train}=0.70$.
@@
We implemented causal trees using the causalTree R package (Athey et al., 2016) and used the default parameters as shown in the example usage on GitHub — that is, split.Rule = "CT", cv.option = "CT", split.Honest = TRUE, cv.Honest = "TRUE", split.Bucket = FALSE, xval = 5, cp = 0, minsize = 20, and propensity = 0.5. As in CDT, we pruned the causal tree using the standard post-pruning procedure, choosing the complexity parameter $\alpha$ which minimizes the cross-validation error.
@@
In the linear and Lasso regressions, we included all main effects $X$ and their interactions with the treatment variable $Z$ as covariates to predict the response $Y$. In the linear regression, we defined a selected subgroup feature to be any covariate whose interaction with the treatment variable yielded a significant p-value ($p<0.05$). In the Lasso regression, we defined a selected subgroup feature to be any covariate whose interaction with the treatment variable received a non-zero coefficient. To estimate the CATE for each unit, we computed the difference between the predicted response when $Z_{i}=1$ and the predicted response when $Z_{i}=0$. The linear regression was implemented using lm(), and the Lasso regression was implemented using glmnet::cv.glmnet() (Friedman et al., 2010) with 5-fold cross-validation and the default hyperparameter grid.
Appendix D Additional Simulation Results
@@
In particular, to complement Figure 4 which examined the subgroup estimation performance under an outcome model including linear covariate effects (i.e., $Y_{i}=Z_{i}\cdot\tau_{i}+X_{i}^{(3)}+X_{i}^{(4)}+\nu_{i}$), we also examined the subgroup estimation performance under the following outcome model, which depends only on the conditional average treatment effect:
@@
Besides this modification in the outcome model, all other simulation settings were kept the same. The results are shown in Figure A2. As seen under the previous outcome model, the distilled methods (Distilled Causal Forest and Distilled Rboost) generally outperformed existing methods in estimating the subgroup structure and the subgroup average treatment effects.
Figure A2: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes in the CATE-only outcome model scenario (rows). CDT with various teacher models (i.e., Distilled Causal Forest and Distilled Rboost) frequently yields the highest $F_{1}$ and number of true positives alongside the lowest number of false positives, threshold RMSEs, and CATE RMSE, demonstrating its effectiveness for accurate subgroup estimation. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@
When examining the distribution of the estimated subgroup thresholds from the tree-based methods in Figure A3, we uncover another benefit of distillation — the distribution of the CDT-estimated subgroup thresholds is much tighter (i.e,. has smaller variance) than for other tree-based subgroup detection methods without distillation (i.e., causal trees and virtual twins). This again reinforces our theoretical understanding of CDT, where we have seen in Example 3.1 that the first-stage learner in CDT acts as a de-noising step, leading to more stable splits (i.e., thresholds) in CDT compared to non-distilled tree-based methods.
Figure A3: Distribution of estimated thresholds for each true subgroup feature (i.e., $X^{(1)}$ and $X^{(2)}$) using different subgroup estimation methods (color) under various treatment effect heterogeneity strengths (columns) and subgroup data-generating processes (subplots). The range of the threshold distributions from CDT (i.e., Distilled Causal Forest and Distilled Rboost) is often smaller than that from causal tree, especially as the treatment effect heterogeneity strength increases. Results are shown for 100 simulation replicates.
@@
Recall thus far that we have been implementing CDT and causal trees with the standard post-pruning procedure using cross-validation to tune the complexity parameter $\alpha$. To investigate the impact of this post-pruning step, we compare the subgroup estimation performance of CDT and causal trees with and without post-pruning in Figure A4. While the performance of causal trees declines drastically without pruning, the performance of CDT with and without pruning is similarly strong, demonstrating CDT’s robustness with respect to this pruning choice.
Figure A4: Performance of pruned and unpruned versions of CDT and causal trees for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). While causal tree without pruning performs substantially worse than pruned causal tree, CDT with various teacher models (i.e., Distilled Causal Forest and Distilled Rboost) performs similarly well with or without pruning. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@
In general, optimally pruning decision trees is a challenging problem and an active area of research (Zhou and Mentch, 2023). To investigate whether the observed difference in subgroup estimation performance between tree-based methods is due to suboptimal pruning or not, we show in Figure A5 the subgroup estimation performance of the tree-based methods, pruned to have a fixed depth of 2 (which is the oracle tree depth given our subgroup DGPs, described in Section 5.1). Even with this oracle pruning, we still observe improvements due to distillation. In particular, compared to causal tree and virtual twins, CDT yields fewer false positives and more accurate estimation of the subgroup ATEs.
Figure A5: Performance of subgroup estimation methods with oracle pruning for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). CDT with various teacher models (i.e., Distilled Causal Forest and Distilled Rboost) remains the most robust and accurate subgroup estimation method, as seen by its high $F_{1}$ and number of true positives alongside low number of false positives, threshold RMSEs, and CATE RMSE. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@
Another choice in the CDT framework is the student model, for which we recommend and have been using a CART decision tree. In previous work (Bargagli-Stoffi et al., 2020; Wan et al., 2023), Rulefit (Friedman and Popescu, 2008) has been proposed to either generate subgroup rules or estimate subgroups with heterogeneous treatment effects. In Figure A8, we compare the performance of Distilled Causal Forest using CART as the student model to Distilled Causal Forest using various instantiations of rulefit as the student model. Namely, we include four different versions of rulefit, implemented using pre::pre() (Fokkema, 2020), with the following hyperparameter settings:

### v2

@@ B.5 Proof of Corollary
[…]
@@
[…]
@@
[…]
@@
[…]
@@ B.7 Proof of Lemmas
$~{}~{}~{}~{}~{}~{}~{}~{}~{}\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}-\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},s^{(a)})\right\}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},\hat{s}^{(a)}_{n})\right\}<0\Big{]}$
@@
$=\Pr\left(\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}-\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]+\delta_{ba}<0\right)$
$=\Pr\left(\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.-\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right]>\delta_{ba}\right)$
$\leq\Pr\left(\left|\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left.\left.\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}_{i};X^{(b)},s^{(b)})\right\}\right]\right|>\delta_{ba}\right)$
@@
$\leq\frac{1}{\delta_{ba}}\mathbb{E}\left(\left|\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.-\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},s^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right]\right|\right)$
$\leq\frac{1}{\delta_{ba}}\left\{\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right|\right]\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.+\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right|\right]\right\}$
@@ Appendix C Method Implementation Details
[…]
@@
[…]
To perform honest estimation of the subgroup ATEs, we set $\pi_{train}=0.70$ for splitting the data into a training and hold-out estimation set. To estimate the individual-level treatment effects $\hat{\tau}(X_{i})$, we used the out-of-bag sample estimates (without repeated cross-fitting) in the Distilled Causal Forest and Distilled BCF, and we used the repeated cross-fitting procedure with $R=50$ repeats in the Distilled Rboost. We also examined other choices of $R$ in Figure A8. From these simulations, we found that $R=50$ provided a good balance between the subgroup estimation accuracy and computational burden. However, more generally, the subgroup estimation results from Distilled Rboost were robust as long as $R$ was sufficiently large enough.
@@
[…]
@@
[…]
@@
[…]
Appendix D Simulations
D.1 Details on Simulation Evaluation Metrics
  1. 1.
  2. 2.
  3. 3.
D.2 Additional Simulation Settings
@@
In particular, to complement Figure 3 which examined the subgroup estimation performance under an outcome model including linear covariate effects (i.e., $Y_{i}=Z_{i}\cdot\tau_{i}+X_{i}^{(3)}+X_{i}^{(4)}+\nu_{i}$), we also examined the subgroup estimation performance under the following outcome model, which depends only on the CATE:
@@
Besides this modification in the outcome model, all other simulation settings were kept the same. The results are shown in Figure A3. As seen under the previous outcome model, the distilled methods (Distilled Causal Forest and Distilled Rboost) generally outperformed existing methods in estimating the subgroup structure and the subgroup ATEs. Moreover, the performance of the linear models for estimating the subgroup ATEs significantly deteriorated under this CATE-only outcome model. This highlights the sensitivity of the linear model approaches to the form of the outcome model. In comparison, CDT tends to work well in both the setting considered here and in Section 5.
Figure A3: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes in the CATE-only outcome model scenario (rows). CDT with various teacher models (i.e., Distilled Causal Forest and Distilled Rboost) frequently yields the highest $F_{1}$ and number of true positives alongside the lowest number of false positives, threshold RMSEs, and subgroup ATE RMSE, demonstrating its effectiveness for accurate subgroup estimation. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@
When examining the distribution of the estimated subgroup thresholds from the tree-based methods in Figure A4, we uncover another benefit of distillation — the distribution of the CDT-estimated subgroup thresholds using Rboost or causal forest is much tighter (i.e,. has smaller variance) than for other tree-based subgroup detection methods without distillation (i.e., causal trees and virtual twins). This again reinforces our theoretical understanding of CDT, where we have seen in Example 3.1 that the first-stage learner in CDT acts as a de-noising step, leading to more stable splits (i.e., thresholds) in CDT compared to non-distilled tree-based methods.
Note that while the thresholds distribution from Distilled BCF is more variable than the non-distilled methods in weak-to-moderate treatment effect heterogeneity regimes, Distilled BCF would have been easily ruled out in these settings according to our teacher model selection procedure using Jaccard SSI (see Appendix E).
Figure A4: Distribution of estimated thresholds for each true subgroup feature (i.e., $X^{(1)}$ and $X^{(2)}$) using different subgroup estimation methods (color) under various treatment effect heterogeneity strengths (columns) and subgroup data-generating processes (subplots). The range of the threshold distributions from Distilled Causal Forest and Distilled Rboost is often smaller than that from causal tree, especially as the treatment effect heterogeneity strength increases. Results are shown for 100 simulation replicates.
@@
Recall thus far that we have been implementing CDT and causal trees with the standard post-pruning procedure using cross-validation to tune the complexity parameter $\alpha$. To investigate the impact of this post-pruning step, we compare the subgroup estimation performance of CDT and causal trees with and without post-pruning in Figure A5. While the performance of causal trees declines drastically without pruning, the performance of CDT using causal forest or Rboost is similarly strong with and without pruning, demonstrating their robustness with respect to this pruning choice. For Distilled BCF, pruning improves the subgroup estimation performance, particularly in the weak-to-moderate treatment effect heterogeneity regimes.
Figure A5: Performance of pruned and unpruned versions of CDT and causal trees for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). While causal tree without pruning performs substantially worse than pruned causal tree, CDT with various teacher models (e.g., Distilled Causal Forest and Distilled Rboost) performs similarly well with or without pruning. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@
[…]
Figure A6: Performance of subgroup estimation methods with oracle pruning for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). CDT with various teacher models remains the most robust and accurate subgroup estimation method, as seen by its high $F_{1}$ and number of true positives alongside low number of false positives, threshold RMSEs, and subgroup ATE RMSE. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@
[…]